OpenStudy (anonymous):

The sum of three consecutive odd integers is 105. What is the first integer? 33 35 37 39

OpenStudy (anonymous):

guess and check

OpenStudy (anonymous):

\(33+35+37=?\)

OpenStudy (anonymous):

Let the first odd integer be : x Other two will be: (x+2) and (x+4) so, x + x +2 + x+ 4 = 105 3x = 99 Now find x from here..

OpenStudy (anonymous):

hey, go it on the first try!

OpenStudy (jiteshmeghwal9):

let the three integers be (x),(x+2),(x+4). then A/q\[(x)+(x+2)+(x+4)=105\]now solve this & gt the value of x. & that will be ur answer:)

OpenStudy (anonymous):

its 33